Circular No. 6856

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 980326 AND GRB 980329 M. S. Briggs, G. Richardson, R. M. Kippen, and P. M. Woods, Uni-

versity of Alabama in Huntsville, report on behalf of the BATSE team: “GRB 980326 (IAUC 6851) was observed with BATSE on Mar. 26.88811 UT as trigger 6660. The event lasted ∼ 5 s and exhibited three narrow pulses. Its peak flux (integrated over 0.5 s) and fluence (50–300 keV) are 8 × 10^{−7} erg cm^{−2} s^{−1} and 1 × 10^{−6} erg cm^{−2}, respectively. GRB 980329 (IAUC 6853) was observed with BATSE on Mar. 29.15600 as trigger 6665; the event was very intense and lasted ∼ 55 s, exhibiting a 10-s-long, highly structured peak. Its peak flux (integrated over 0.5 s) and fluence (50–300 keV) are 8 × 10^{−6} erg cm^{−2} s^{−1} and 5 × 10^{−5} erg cm^{−2}, respectively. The BATSE locations are consistent with the locations of the reported optical transient for GRB 980326 (IAUC 6852) and the SAX/NFI x-ray counterpart for GRB 980329 (IAUC 6854). Location maps can be found at

http://www.batse.msfc.nasa.gov/~kippen/batsebr.”

S. A. Ilovaisky and C. Chevalier, Observatoire de Haute-Provence (OHP), write: “Two 15-min R-band CCD exposures of the SAX NRI error box for the GRB 980329 γ-ray burst (IAUC 6853, 6854) were taken by T. Bremnes, University of Basel, with the 1.2-m OHP telescope on Mar. 31.87 UT (65 hr after the burst). More than 15 objects are seen down to mag 21 within the 1′ error circle. Comparison with a film copy of the second Palomar Sky Survey red plate for this region reveals one object within the NFI error circle that is present in our frames and not on the Palomar plate. This object, for which we derive a preliminary magnitude of R = 20.6, is located near a bright (R = 15.8) star, and has coordinates α = 7^{h}02^{m}38^{s}.8, δ = +38^{o}50′27′′ (equinox 2000.0), as derived using the Digitized Sky Survey and the USNO-A1 catalogue. This is 30′′ east of the center of the 1′ radius error box. Our average R-band image is available at

http://www.obs-hp.fr/~ilovaisky/grb980329.html. Further imaging

of this field is needed.”

XTE J0421+560 W. Paciesas, University of Alabama in Huntsville; and G. Fishman,

Marshall Space Flight Center, NASA report for the CGRO/BATSE team: “This new x-ray transient (IAUC 6855) was also detected by BATSE, beginning on Mar. 31.91 UT, at a flux of 470 ± 50 mCrab (20–100 keV). The source was not significantly detected during Mar. 31.03–31.84, with an upper limit of 55 mCrab (3σ).”

1998 April 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green